# Secluded Dark Composites and Remnant Binding Fields

### v1

@@ 1 Introduction
Dark matter’s identity and interactions remain unsolved mysteries in the modern scientific era. Amongst the many possible dark matter candidates, composite dark matter has some unique formation pathways and interactions, which are still being explored[1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. For example, it was recently shown that a light scalar field, which produces an attractive potential that binds together fermionic composites, can produce multiple low-energy recoils in underground dark matter search experiments [35] and even Bremmstrahlung radiation from Standard Model nuclei accelerated in the dark matter composite’s interior [34]. This and other novel searches for composite dark matter (e.g. [15, 19, 20, 21, 22, 23, 39, 40, 41, 42, 38, 43]) motivate a deeper understanding of its origins.
@@
The rest of this paper is organized as follows. In Section II, we study the freeze-out dynamics of secluded composite dark matter, focusing on the annihilation process $\chi+\bar{\chi}\to\varphi+\varphi$, the subsequent formation of bound states, and the energy density evolution of the remnant scalar field $\varphi$ during composite assembly. We examine the contributions of $\varphi$ to cosmological observables including the effective number of relativistic species ($N_{\mathrm{eff}}$), and detail constraints from Planck data and projected CMB-S4 sensitivities. In Section III, we introduce a UV thermal coupling which ensures DM-SM thermal equilibrium during the early universe and sets the DM cosmological abundance. Further, we introduce a Higgs portal interaction to connect the dark sector to the Standard Model. This interaction allows us to derive bounds on the mass and coupling of the Higgs portal scalar. We then use these results to place constraints on direct detection prospects for this model, considering both astrophysical and experimental limits. Section IV concludes with a summary of our results and a discussion of future directions. Throughout this work, we use natural units where $\hbar=c=k_{B}=1$.
@@ 2.1 Secluded asymmetric dark sector abundance
In the upper left panel of Figure (1), we evaluate the energy densities of $\chi$, $\varphi$ and the Standard Model as functions of the Standard Model radiation temperature, and show the cosmological abundances in the case that asymmetric $\chi$s do not form composites.
@@
Figure 1: Schematic cosmological energy densities, matching Boltzmann equation outputs, are shown as a function of the Standard Model radiation temperature $T$, for a secluded dark sector which can form composites out of equilibrium with the Standard Model. The secluded dark matter field $\chi$ annihilates solely to scalar field $\varphi$, leaving behind a residual dark matter density “$\chi,$ res”. The energy densities of the symmetric dark matter, asymmetric dark matter residual, Standard Model radiation and light scalar $\varphi$ components are represented by the blue, red, orange, and green curves respectively. In the top left panel, we show the case where no DM composite forms for comparison. In all other panels, we set $m_{X}$ = 10 GeV, $\alpha$ = 0.1, and vary the mass of the light scalar $m_{\varphi}$ as indicated. The asymmetric residual population of $\chi$s form composites at a lower temperature, during which process the binding energy of the composites is emitted as $\varphi$. The initial residual dark matter density “$\chi$, res” is set so that matter-radiation equality is obtained at $T=0.8$ eV. The initial $\varphi$ and $\chi$ abundance are set to reflect having previously been in thermal equilibrium with the SM at high temperature.
@@ 2.2 Bound state formation
We now turn to the structure and size of the composites. At the temperature of composite assembly, the expected number of constituents for a strongly-bound ($BE/N\simeq m_{\chi}$) composite is given by [35],
$N$ $=\left(\frac{20\sqrt{g_{\mathrm{ca}}^{*}}T_{\mathrm{m-r}}T_{\mathrm{ca}}^{3/2}M_{\mathrm{pl}}}{\bar{m}_{\chi}^{7/2}}\right)^{6/5}$ (22)
$\simeq 10^{14}\left(\frac{g_{\mathrm{ca}}^{*}}{10}\right)^{3/5}\left(\frac{T_{\mathrm{ca}}}{1\,\mathrm{GeV}}\right)^{9/5}\left(\frac{1\,\mathrm{GeV}}{\bar{m}_{\chi}}\right)^{21/5},$ (23)
@@ 3.2 Higgs Portal
Figure 6: Exclusion regions for the mixing angle $\theta$ as a function of the scalar mass $m_{\varphi}$ in a Higgs portal scalar model. Shaded regions represent constraints from stellar observations (red) and cosmology (green) (both from [59]), and fifth-force experiments (blue). The fifth-force bounds were derived by recasting limits on ${m_{\varphi},\alpha_{\mathrm{grav}}}$ in [60, 61, 62, 63] in terms of $\sin\theta$ via Equation (38). The orange dotted lines correspond to the saturation of the condition $\lambda_{h\varphi}<4\times 10^{-8}$, above these lines (for indicated values of $\lambda_{\phi}$) the $\varphi$ abundance from $h\to\varphi\varphi$ exceeds the dark matter abundance. The multiplicity of these lines arises because $\theta$ and $\lambda_{h\varphi}$ are degenerate without a specified scalar self-coupling $\lambda_{\varphi}$: using $v_{\varphi}=m_{\varphi}/\sqrt{2\lambda_{\varphi}}$ and assuming $m_{\varphi}\ll m_{h}$, the mixing angle scales as $\sin\theta\approx\lambda_{h\varphi}v_{h}m_{\varphi}/(m_{h}^{2}\sqrt{2\lambda_{\varphi}})$.
@@
Next we examine the limits on a Higgs-portal dark scalar from equivalence principle ($aka$ fifth-force) tests for the Yukawa potential
@@
where this potential is normalized according to conventions used in torsion pendulum fifth-force experiments; $G$ is Newton’s constant and $m_{N}$ is the nucleon mass [constraint1, constraint2]. Matching this to the Higgs portal Yukawa term in Equation (34) we find

### v2

@@ 1 Introduction
Dark matter’s identity and interactions remain unsolved mysteries in the modern scientific era. Amongst the many possible dark matter candidates, composite dark matter has some unique formation pathways and interactions, which are still being explored[68, 34, 25, 40, 30, 61, 9, 8, 54, 59, 56, 57, 55, 27, 53, 67, 46, 47, 16, 44, 15, 52, 42, 22, 11, 29, 10, 19, 13, 21, 43, 45, 66, 3, 4, 35, 28, 60, 69]. For example, it was recently shown that a light scalar field, which produces an attractive potential that binds together fermionic composites, can produce multiple low-energy recoils in underground dark matter search experiments [4] and even Bremsstrahlung radiation from Standard Model nuclei accelerated in the dark matter composite’s interior [3]. This and other novel searches for composite dark matter (e.g. [53, 16, 44, 15, 52, 42, 12, 23, 24, 2, 60, 18]) motivate a deeper understanding of its origins.
@@
The rest of this paper is organized as follows. In Section 2, we study the freeze-out dynamics of secluded composite dark matter, focusing on the annihilation process $\chi+\bar{\chi}\to\varphi+\varphi$, the subsequent formation of bound states, and the energy density evolution of the remnant scalar field $\varphi$ during composite assembly. We examine the contributions of $\varphi$ to cosmological observables including the effective number of relativistic species ($N_{\mathrm{eff}}$), and detail constraints from Planck data and projected CMB-S4 sensitivities. In Section 3, we introduce a UV thermal coupling which ensures DM-SM thermal equilibrium during the early universe and sets the DM cosmological abundance. Further, we introduce a Higgs portal interaction to connect the dark sector to the Standard Model. This interaction allows us to derive bounds on the mass and coupling of the Higgs portal scalar. We then use these results to place constraints on direct detection prospects for this model, considering both astrophysical and experimental limits. Section 4 concludes with a summary of our results and a discussion of future directions. Throughout this work, we use natural units where $\hbar=c=k_{B}=1$.
@@ 2.1 Secluded asymmetric dark sector abundance
[…]
@@
[…]
@@ 2.2 Bound state formation
In Figure 3, we show the results of this calculation explicitly in the case where $\alpha$ = 0.1 and $N$ = 10.
Figure 3: Ratio of formation rate to decay rate as a function of constituent mass $m_{\chi}$ for $\alpha$ = 0.1 and $N$ = 10, where $\Gamma_{\mathrm{form}}=H(T_{\mathrm{form}})$, and the decay rate is computed using Eq. (14). We see that $\Gamma_{form}/\Gamma_{decay}\ll 1$, justifying the assumption that the composite decays rapidly into its ground state for composites considered in this work.
We now turn to the structure and size of the composites. At the temperature of composite assembly, the expected number of constituents for a strongly-bound ($BE/N\simeq m_{\chi}$) composite is given by equating the N-composite formation rate with the Hubble rate, which yields the estimate [4, 46],
$N$ $=\left(\frac{20\sqrt{g_{\mathrm{ca}}^{*}}T_{\mathrm{m-r}}T_{\mathrm{ca}}^{3/2}M_{\mathrm{pl}}}{\bar{m}_{\chi}^{7/2}}\right)^{6/5}\simeq 10^{14}\left(\frac{g_{\mathrm{ca}}^{*}}{10}\right)^{3/5}\left(\frac{T_{\mathrm{ca}}}{1\,\mathrm{GeV}}\right)^{9/5}\left(\frac{1\,\mathrm{GeV}}{\bar{m}_{\chi}}\right)^{21/5},$ (22)
@@
Figure 5: The same as Figure 4, but for heavier $m_{\varphi}=10,\,100$ eV. Here we show the relative $\varphi-SM$ density, $\rho_{\varphi}/\rho_{SM}$ at matter-radiation ($i.e.$ $\chi$-SM) equality. Here again it is assumed that the N-body composite decays into its ground state immediately after forming. The gridded dark grey region is where $\alpha$ is below the minimum value required to form composites, and the solid grey region is where the composite would form and decay after matter-radiation equality. The region below the orange-red line is where $\rho_{\varphi}/\rho_{SM}$ $\geq$ 1, which is excluded by requiring matter-radiation equality occur at $T\sim$ eV. The thicker lines and darker grey region represent results for $T_{form}$ = BE/10, while the thinner lines and lighter grey region are for $T_{form}$ = BE/30; this range of formation temperatures corresponds to the uncertainty arising from Saha-type corrections to composite synthesis.
@@ 3.2 Higgs Portal
Figure 7: Exclusion regions for the mixing angle $\theta$ as a function of the scalar mass $m_{\varphi}$ in a Higgs portal scalar model. Shaded regions represent constraints from stellar observations (red) and cosmology (green) (both from [49]), and fifth-force experiments (blue). The stellar cooling bounds are comparable to the white dwarf bounds obtained recently in [14]. The fifth-force bounds were derived by recasting limits on $m_{\varphi},\alpha_{\mathrm{grav}}$} in [5, 26, 65, 70] in terms of $\sin\theta$ via Equation (37). The orange dotted lines correspond to the saturation of the condition $\lambda_{h\varphi}<4\times 10^{-8}$, above these lines (for indicated values of $\lambda_{\phi}$) the $\varphi$ abundance from $h\to\varphi\varphi$ exceeds the dark matter abundance. The multiplicity of these lines arises because $\theta$ and $\lambda_{h\varphi}$ are degenerate without a specified scalar self-coupling $\lambda_{\varphi}$: using $v_{\varphi}=m_{\varphi}/\sqrt{2\lambda_{\varphi}}$ and assuming $m_{\varphi}\ll m_{h}$, the mixing angle scales as $\sin\theta\approx\lambda_{h\varphi}v_{h}m_{\varphi}/(m_{h}^{2}\sqrt{2\lambda_{\varphi}})$. We note that the cosmological constraints shown in green originate from electromagnetic energy injection in the post-recombination Universe: decays occurring at timescales between approximately $10^{6}$ s and $10^{9}$ s lead to $\mu$ and $y$-type spectral distortions in the CMB, while later decays produce photons that survive to presence day and either contribute to or distort the extragalactic background light (EBL) spectrum [38, 48]. These bounds are effective up to scalar masses of order $m_{\varphi}\sim 10^{6}\rm eV$. Further breakdown of cosmological constraints on Higgs Portal scalars is found in [39]. For heavier scalars and larger mixing parameter (not shown in figure), additional constraints arise from Big Bang Nucleosynthesis, where late-time decays can disrupt nuclear process or inject entropy, thereby affecting the baryon-photon ratio. These limits exclude mixing angles up to $\sin\theta\sim 10^{-5}$ and extend to masses as large as $m_{\varphi}\sim\rm 200$ MeV. For even larger mixing angles, $\sin\theta\gtrsim 10^{-4}$, collider searches - in particular Kaon decays - impose additional exclusions [41]. As a result, there remains another narrow window—less than one order of magnitude in $\sin\theta$—between the upper edge of the BBN bounds and the lower edge of the collider bounds where the model is still viable.
@@
Next we examine the limits on a Higgs-portal dark scalar from equivalence principle (akaaka fifth-force) tests for the Yukawa potential
@@
where this potential is normalized according to conventions used in torsion pendulum fifth-force experiments; $G$ is Newton’s constant and $m_{N}$ is the nucleon mass. Matching this to the Higgs portal Yukawa term in Equation (33) we find